A projectile is fired into the air from the top of 200-m cliff above a valley. Its initial velocity is 60m/s at 60° above the horizontal. Where does the projectile land? Show the given,unknown,and solution.

1
Expert's answer
2022-01-10T09:10:08-0500

Given: "y_0=200\ m" is the initial height, "v_0=60\ \dfrac{m}{s}" is the initial velocity, "\theta=60^{\circ}" is the launch angle.

Unknown: "x", the range of the projectile.

Solution:

Let's first find the total time in air from the kinematic equation:


"y=y_0+v_{0y}t+\dfrac{1}{2}gt^2,""0=200+60t\times sin60^{\circ}+\dfrac{1}{2}\times(-9.8)t^2,""4.9t^2-52t-200=0."

This quadratic equation has two roots: "t_1=13.6\ s", "t_2=-2.99\ s". Since time can't be negative the correct answer is "t=13.61\ s".

Finally, we can find the range of the projectile as follows:


"x=v_{0x}t=v_0tcos\theta,""x=60\ \dfrac{m}{s}\times 13.6\ s\times cos60^{\circ}=408\ m."